### v1

# Spatial Proportional Hazards Model with Differential Regularization

Spatial Proportional Hazards Model with Differential Regularization
Lorenzo Tedesco
1 Introduction
Survival analysis has long been an essential tool across various fields, including medicine, engineering, and social sciences, for analyzing time-to-event data. In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards model, first introduced by Cox (1972). This model assesses the effect of covariates on the hazard, or risk, of an event occurring.
In recent years, survival analysis has made significant strides in modeling different forms of dependence in survival data. This paper explores a scenario where the occurrence of an event is associated with its geographic location, introducing a spatial dependence into the data. Specifically, an event’s occurrence may be influenced by spatial factors related to the location where the event happens. We refer to the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards model.
Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions and our method is able to deal with both the framework. A comprehensive literature review, including Bayesian approaches, can be found in Hanson and Zhou (2014). Our interest lies in a frequentist approach to the spatial Proportional Hazards model, for which various models and methods are now available to facilitate inference. Examples include the semiparametric estimating equations method proposed by Li and Lin (2006) and the composite likelihood method Paik and Ying (2012). The main issue with these proposals is that the dependence structure does not consider the shape of the domain but only the distance among the locations. This can be problematic when the boundary of the domain plays a crucial role in the spatial dependence. In this paper, we propose a novel approach based on recent developments in differential regularization, as discussed in Wilhelm and Sangalli (2016). One of the main features of this approach is to be able to deal with domains having complex shapes, characterized for instance by strong concavities or holes, and where the shape of the domain influences the behavior of the phenomenon.
We demonstrate that the methodology introduced in Wilhelm and Sangalli (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data. Specifically, we optimize a penalized log-likelihood function that includes a roughness penalty term based on a differential quantity of the spatial field, calculated over the domain of interest. This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli, 2016). Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field, particularly in domains with complex geometries, where triangulation effectively captures the intricacies of the area.
The paper is organized as follows: In Section 2, we introduce the Spatial Proportional Hazards model. Section 3 discusses the functional version of the PIRLS, tailored to our specific case of interest. In Section 4, we describe the numerical implementation of the method using the finite element method. Section 5 specifies the model version adapted for areal data. Section 6 is dedicated to simulation studies, and Section 7 presents an empirical application of the method. Finally, Section 8 provides the conclusions. All technical details and proofs are deferred to the Appendix.
2 Model
We consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with a boundary $\partial\Omega$ that is regular and belongs to $C^{2}(\mathbb{R}^{2})$. Let $p_{1},\ldots,p_{n}\in\Omega$ denote $n$ points, where $p_{i}=(p_{1i},p_{2i})$. Each point $p_{i}$ is associated with a time-to-event random variable $T_{i}$. Due to censoring, $T_{i}$ cannot be observed directly; instead, we obtain a realization of $(Y_{i},\Delta_{i})$, denoted by $(y_{i},\delta_{i})$, where $Y_{i}=\min(T_{i},C_{i})$, $\Delta_{i}=I(T_{i}\leq C_{i})$, $C_{i}$ is a censoring random variable, and $I(\cdot)$ is the indicator function. Additionally, we observe a covariate vector $x_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$. We assume that $T_{i}$ is independent of $C_{i}$ conditional on $x_{i}$. We use a proportional hazard model for the random variables $\{T_{i}\}_{i=1}^{n}$. Specifically, let $\lambda_{i}(t\mid x)$ denote the conditional hazard function of $T_{i}$ given $x_{i}=x$. We assume that
$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(x^{\top}\beta+f(p_{i})),$ (1)
where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function over $\Omega$ belonging to a suitable functional space $\mathcal{F}$. Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $p^{*}\in\Omega$ it holds $f(p^{*})=0$.
We propose to estimate the regression coefficients $\beta\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
[…]
where $\mathcal{L}(\beta,f)$ is the standard partial likelihood defined as
$\mathcal{L}(\beta,f)=\sum_{i:\delta_{i}=1}\left(x_{i}^{\top}\beta+f(p_{i})-\log\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\beta+f(p_{j})\right)\right).$ (3)
Here, $\mu$ is a positive smoothing parameter, and the Laplacian $\Delta f=\frac{\partial^{2}f}{\partial p_{1}^{2}}+\frac{\partial^{2}f}{\partial p_{2}^{2}}$ measures the local curvature of the field $f$. A higher $\mu$ controls the wiggliness of $f$, while a smaller $\mu$ allows for more flexibility in $f$.
We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al. (2013). The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\beta}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$. As in Wilhelm and Sangalli (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al. (1986) for maximizing a penalized log-likelihood functional in generalized additive models. This is because the penalized log-likelihood in (2) involves a function parameter, the spatial field $f$, and the maximization occurs over the space $\mathbb{R}^{q}\times\mathcal{F}$, where $\mathcal{F}$ is infinite-dimensional. Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli (2016) to find an approximate solution to the estimation problem.
3 Functional version of the PIRLS algorithm
We consider the more general penalized log-likelihood function given by:
$\mathcal{L}_{p}(\beta,f)=\mathcal{L}(\beta,f)-\frac{\mu}{2}m(f,f),$ (4)
where $\mathcal{L}$ denotes the log-likelihood and $m(\cdot,\cdot):\mathcal{F}\times\mathcal{F}\to\mathbb{R}$ is a bilinear, symmetric, and semi-positive definite form. This formulation enables us to extend the PIRLS algorithm to accommodate any functional roughness penalty of this general quadratic form.
For that, define the following quantities:
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
$r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f),$
$l_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)},$
$v_{i}(\beta,f)$ $=\theta_{i}(\beta,f)/l_{i}(\beta,f).$
First, we show that the problem of maximizing 4 is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant. As, in fact, $l$ and $V$ also depend on $(\beta,f)$, this suggests an iterative scheme for the solution of the estimation problem. Let $(\beta^{(k)},f_{n}^{(k)})$ be the value of $(\beta,f)$ after $k$ iterations of the algorithm, and let $\theta^{(k)}$ the estimate of the vector $\theta$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\theta^{(k)})$. Define also $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X^{\top}\beta^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$. The first order approximation of $\theta(\beta,f)$ in the neighborhood of the current value of $\theta^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X^{\top}\beta-f)\right\|^{2}+\mu m(f,f).$ (5)
We can consider the following iterative scheme:
  1. 1.
    Compute $z^{(k)}$ and $W^{(k)}$ based on $(\beta^{(k)},f_{n}^{(k)})$.
  2. 2.
    Next, find $\beta$ and $f_{n}$ that minimize (5).
  3. 3.
    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\beta}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
  4. 4.
    Define $\beta^{(k+1)}=\hat{\beta}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
Here $f_{n}^{(k)}(p^{*})$ is the estimation of $f$ at step $k$ at $p^{*}$. The standardization of $f_{n}^{(k)}$ in step 4 ensures that $f_{n}(p^{*})=0$ and that (1) is satisfied. We can use a stopping criteria based on a sufficiently small variation of two successive values of (5) or a maximum number of iterations.
4 Numerical Implementation using the Finite Element Method
[…]
$\tilde{J}_{\lambda}(\beta,f)=\|W^{1/2}(z-X\beta-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
For simplicity, we omit the dependence on the iteration counter $k$ in this notation and throughout the following sections. Let us now consider what type of space $\mathcal{F}$ is suitable for the current problem. To do so, we introduce the Sobolev space $H^{m}(\Omega)$, which is a Hilbert space comprising all functions in $L^{2}(\Omega)$ along with all their distributional derivatives up to order $m$. Since the roughness penalty term $\int_{\Omega}(\Delta f)^{2}$ must be well-defined, it follows that $\mathcal{F}\subset H^{2}(\Omega)$. According to the Sobolev embedding theorem, $H^{2}(\Omega)\subset C^{0}(\Omega)$. Therefore, any function $f\in H^{2}(\Omega)$ is continuous and can be evaluated pointwise, allowing us to compute the vector $f_{n}$ in the partial likelihood term. To ensure the uniqueness of the minimizer in (6), appropriate boundary conditions must be imposed. These conditions specify the desired behavior of the estimated function $f$ at the boundaries of the domain of interest. Typically, constraints can be set on the value of $f$ at the boundary $\partial\Omega$, i.e., $f|_{\partial\Omega}=\gamma_{D}$ (Dirichlet boundary conditions), or on the flux of the function through the boundary, i.e., $\partial_{n}f|_{\partial\Omega}=(\nabla f)^{\top}\mathbf{n}=\gamma_{N}$ (Neumann boundary conditions), where $\mathbf{n}$ is the outward-pointing normal unit vector to the boundary, and $\nabla f=(\partial f/\partial p_{1},\partial f/\partial p_{2})^{\top}$ represents the gradient of $f$. If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are termed homogeneous. It is also possible to apply different boundary conditions to various parts of the boundary, forming a partition of $\partial\Omega$. To ensure the uniqueness of the minimization problem (4.1), we consider the space:
$\mathcal{F}=H^{2}_{n}=\left\{f\in H^{2}\mid(\nabla f)^{\top}\mathbf{n}=0\text{on }\partial\Omega\right\}.$
For more details on general boundary conditions, see Azzimonti et al. (2014).
4.1 Characterization of the Solution to the Penalized Partial Likelihood Problem
We assume that the design matrix $XX^{\top}$ is of full rank and that the weight matrix $W$ contains strictly positive entries. Define $H=X(X^{\top}WX)^{-1}X^{\top}W$ and $Q=I-H$, where $I$ is an identity matrix of the appropriate size. For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $u_{n}=(u(p_{1}),\ldots,u(p_{n}))$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{p_{i}\}_{i=1}^{n}$. Let $\tilde{\beta}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\beta,f)$ in (4). Under these conditions, the following Proposition describes the minimizers $\tilde{\beta}$ and $\tilde{f}$ of the functional (6).
Proposition 1.
There exists a unique pair $(\tilde{\beta},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
  • $\tilde{\beta}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{f}_{n}=(\tilde{f}(p_{1}),\ldots,\tilde{f}(p_{n}))^{\top}$,
  • $\tilde{f}$ satisfies:
    $u_{n}^{\top}Q\tilde{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=u_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
Using Proposition 1 and the functional version of the PIRLS algorithm from Section 3, we obtain a characterization of the maximum penalized log-likelihood within the functional space $H^{2}_{n}(\Omega)$.
4.2 Solution to the Penalized Partial Likelihood
Here, we outline the methodology for minimizing $\tilde{J}_{\lambda}(\beta,f)$ with respect to both $\beta$ and $f$. As indicated by Proposition 1, computing $\tilde{\beta}$ is straightforward. The key challenge is to find $\tilde{f}$ that satisfies (7). To address this, we introduce the space:
$H^{1}_{n}=\left\{f\in H^{1}\mid(\nabla f)^{\top}\mathbf{n}=0\text{ on }\partial\Omega\right\}.$
According to Wilhelm and Sangalli (2016); Sangalli et al. (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
$\begin{cases}u_{n}^{\top}Q\tilde{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=u_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
for each $(u,v)\in H^{1}_{n}\times H^{1}_{n}(\Omega)$. This formulation requires less regularity on the functions involved compared to formulation (6) defined in $H^{2}_{n}(\Omega)$. In the following section, we discuss how the finite element method can be employed to construct a finite-dimensional subspace of $H^{1}_{n}(\Omega)$, enabling the computation of an approximate solution to (7) within this space.
4.3 Finite Elements
The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni, 2009).
To build a finite element space, we begin by dividing the domain of interest $\Omega$ into smaller subdomains. Common domain partitions include triangular meshes. We consider a regular triangulation $\mathcal{T}$ of $\Omega$, where adjacent triangles share either a vertex or an edge. The domain $\Omega$ is thus approximated by $\Omega_{\mathcal{T}}$, which is the union of all triangles. Consequently, the boundary $\partial\Omega$ is approximated by a polygon (or multiple polygons for domains with interior holes). This triangulation effectively captures the complex geometry of the domain, including significant concavities such as rivers and detailed structures like bridges that connect different parts of the city.
From this triangulation, locally supported polynomial functions are defined over the triangles, forming a set of basis functions $\psi_{1},\ldots,\psi_{K}$ that span a finite-dimensional subspace $\mathcal{F}_{K}$ of $H^{1}_{n}$. Linear finite elements are created using a basis system where each basis function $\psi_{i}$ is associated with a vertex $\xi_{i}$ (for $i=1,\ldots,K$) of the triangulation $\mathcal{T}$. Each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta. Figure 1 provides an example of a linear finite element basis function on a planar mesh, highlighting its localized support.
Let $\psi=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$. Any function $h$ in the finite element space $\mathcal{F}_{K}$ can be expressed as a linear combination of the basis functions $\psi_{1},\ldots,\psi_{K}$. Let $h=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $h$, such that
$h(\cdot)=\sum_{j=1}^{K}h_{j}\psi_{j}(\cdot)=\mathbf{h}^{\top}\psi(\cdot).$
Note that
$h(\xi_{i})=\sum_{j=1}^{K}h_{j}\psi_{j}(\xi_{i})=\sum_{j=1}^{K}h_{j}\delta_{ij}=h_{i},$
therefore,
$\mathbf{h}=(h(\xi_{1}),\ldots,h(\xi_{K})),$
which shows that each function $h\in\mathcal{F}_{K}$ is completely determined by its values at the mesh nodes.
Figure 1: Example of a linear finite element basis function on a planar mesh. The function is locally supported, taking the value one at a specific vertex and zero at all other vertices of the mesh.
4.4 Numerical solution to the Penalized Partial Likelihood
The functions and integrals in equation (8) can be approximated using elements from the finite element space $\mathcal{F}_{K}$. As a result, the problem (8) is approximated by its discrete version: find $(\tilde{f},\tilde{h})\in\mathcal{F}_{K}\times\mathcal{F}_{K}$ such that (8) holds for all $(u,v)\in\mathcal{F}_{K}\times\mathcal{F}_{K}$, where the integrals are now computed over the triangulation $\Omega_{\mathcal{T}}$. Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\mathbf{p}_{1},\ldots,\mathbf{p}_{n}$,
$\Psi=\begin{bmatrix}\psi^{\top}(\mathbf{p}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{p}_{n})\end{bmatrix}$ (9)
and consider the $K\times K$ matrices
$\mathbf{R}_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad\mathbf{R}_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
Using this notation, for functions $\tilde{f},\tilde{h},u,v\in\mathcal{F}_{K}$, the integrals in (8) can be written as:
$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}\mathbf{R}_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}\mathbf{R}_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}\mathbf{R}_{0}\mathbf{v},$
where $\tilde{\mathbf{f}},\tilde{\mathbf{h}},\mathbf{u}$, and $\mathbf{v}$ are the coefficient vectors of the basis expansions of $\tilde{f},\tilde{h},u$, and $v$, respectively. Consequently, the discrete version of problem (8) reduces to solving a linear system, as formulated in the following proposition.
Proposition 2.
The discrete form of (8) is given by the system
$\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\Psi&\lambda\mathbf{R}_{1}\\ \lambda\mathbf{R}_{1}&\lambda\mathbf{R}_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\mathbf{z}\\ 0\end{bmatrix},$ (10)
which has a unique solution pair $(\tilde{\mathbf{f}},\tilde{\mathbf{h}})$ corresponding to the coefficients of the basis expansions of $\tilde{f}$ and $\tilde{h}$.
The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $\mathbf{R}_{0}$ and $(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1})$.
Let $\mathbf{P}=\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1}$ and $\mathbf{S}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}$. Using the functional form of the PIRLS algorithm, and leveraging Propositions 1 and 2, we obtain the following expressions for the maximizers $\hat{\bm{\beta}}$ and $\hat{f}$ of the penalized partial likelihood (2):
$\hat{\bm{\beta}}=(\mathbf{X}^{\top}\mathbf{W}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{W}(\mathbf{I}-\mathbf{S})\mathbf{z},$
with $\hat{f}$ identified by the vector
$\hat{\mathbf{f}}=(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z},$
and the vector of evaluations of $\hat{f}$ at the $n$ data locations given by
$\hat{\mathbf{f}}_{n}=\Psi\hat{\mathbf{f}}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z}=\mathbf{S}\mathbf{z},$
where the vector of pseudo-data $\mathbf{z}$, and the matrices $\mathbf{W},\mathbf{Q}$, and $\mathbf{S}$ are those obtained upon convergence of the PIRLS algorithm.
The positive definite matrix $\mathbf{P}$ represents the discretization of the penalty term in (2) and (7). Due to the variational formulation (8) of the estimation problem, this penalty matrix does not require the computation of second-order derivatives. Azzimonti et al. (2014) demonstrates that in the finite element space used for discretization, $\mathbf{P}$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
4.5 Selecting the Parameter $\mu$ Using Cross-Validation
When it comes to choosing the parameter $\mu$, one straightforward approach is a visual inspection. This method allows the researcher to select a value that achieves the desired smoothness of the estimated spatial field. However, to avoid subjective decision-making, we propose a more objective, numerical method. Following the approach of Cygu et al. (2021), we can utilize cross-validation to determine the optimal value of $\mu$. Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al., 1996).
To compute the CV-PLD for each $\mu$, we employ $k$-fold cross-validation. This process involves splitting the training data into $k$ subsets (folds), training the model on $k-1$ folds, and validating it on the remaining fold. This procedure is repeated $k$ times, each time with a different fold serving as the validation set. We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny, 2019). The CV-PLD is defined as:
[…]
where $\mathcal{L}(\hat{\beta}_{-k}(\mu))$ represents the log partial likelihood evaluated at $\hat{\beta}_{-k}$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\beta}_{-k}(\mu))$ is the log partial likelihood evaluated at $\hat{\beta}_{-k}$ on the retained data (all but the excluded fold). The $\hat{\beta}_{-k}$ values denote the penalized estimates derived from the retained data. The optimal value of $\mu$ is the one that minimizes (11). Notably, equation (11) typically provides superior results compared to simply evaluating the partial likelihood on the held-out fold (a method sometimes referred to as the basic approach), as the likelihood of any given observation is influenced by the other elements in the risk set.
The alternative cross-validation metric, the CV-C-index, is based on Harrell’s concordance index for Cox models. This metric, known as the cross-validated C-index, calculates the probability that, for a random pair of individuals, the predicted survival times are correctly ordered according to their actual survival times. This method is similar to the implementation available in the survival package (Therneau et al., 2015).
5 Model Version for Areal data
//TODO
6 Simulation Results
To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay (2002); Wood et al. (2008). We use the spatial field $f(x,y)$, depicted in Figure 2(a) and defined as follows:
$f(x,y)=\begin{cases}-\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y-r)^{2}\leq(r-r_{0})^{2}\text{ and }x\geq 3,\\ -\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,r_{0}\leq y\leq 2r-r_{0},\\ -\frac{1}{8}\left(\arctan\left(-\frac{y}{x}\right)r+\left(\sqrt{x^{2}+y^{2}}-r\right)^{2}+10\right)+c&\text{if }r_{0}^{2}\leq x^{2}+y^{2}\leq(2r-r_{0})^{2}\text{ and }x<0,\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,-(2r-r_{0})\leq y\leq-r_{0},\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y+r)^{2}\leq(r-r_{0})^{2}\text{ and }x>3,\\ 0&\text{otherwise},\end{cases}$
where $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$, $r=0.5$ and $r_{0}=0.1$.
We randomly sample $n=100,200,300$ points, denoted by $\{p_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b). Figure 2(c) shows the corresponding Delaunay triangulation. Each sampled point is associated with an observation $x_{i}$ of a random variable $X_{i}$, which follows a standard normal distribution. Additionally, we observe $(y_{i},\delta_{i})$, where $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=I(T_{i}\leq C_{i})$. The variable $T_{i}$ given $X_{i}=x$ follows a proportional hazards model with baseline hazard $\lambda_{0}(t)=1$ and $\beta_{0}=0.2$, as defined in equation (1), with the form $\lambda(t|x)=\lambda_{0}(t)\exp(x\beta_{0}+f(p_{i}))$. To introduce censoring, we set $C_{i}=1$ for each $i$, resulting in approximately 20% censored observations. We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ effects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$. The finite elements method is applied using a basis of order $1$ constructed on the Toggle the table of contents Delaunay triangulation over the sampled locations.
(a) Spatial field $f(x,y)$ over the horseshoe domain
(b) Sampled points over the horseshoe domain
(c) Mesh based on the sampled points
Figure 2: Figures illustrating various aspects of the spatial field and sampling.
We conduct $N=100$ simulations for each combination of $n$ and $\mu$. And example of the estimated spatial field in Figure 3. This figure compares the true values (Figure 3(a)) with the estimated values (Figure 3(b)), and reports the pointwise error. In table
(a) True Values
(b) Estimated Values
(c) Error Values
Figure 3: Comparison of true values, estimated values, and error values.
$n$ $\mu$ MSE of $f$ (sd) Bias of $\beta$ (sd)
100 0.5 0.168 (0.112) -0.008 (0.130)
1 0.168 (0.118) -0.007 (0.130)
3 0.177 (0.140) -0.006 (0.130)
5 0.186 (0.165) -0.006 (0.130)
10 0.190 (0.188) -0.007 (0.129)
200 0.5 0.157 (0.119) 0.008 (0.096)
1 0.160 (0.129) 0.007 (0.096)
3 0.168 (0.150) 0.007 (0.095)
5 0.180 (0.169) 0.007 (0.094)
10 0.185 (0.179) 0.008 (0.093)
300 0.5 0.141 (0.105) -0.003 (0.077)
1 0.142 (0.115) -0.002 (0.077)
3 0.143 (0.129) 0.000 (0.077)
5 0.144 (0.136) 0.000 (0.077)
10 0.148 (0.149) 0.000 (0.077)
Table 1: Simulation results for varying sample sizes ($n$) and smoothing parameters ($\mu$), showing the mean squared error (MSE) of the spatial field estimator $f$ and the bias of the parameter $\beta$. Standard deviations (sd) are reported in parentheses for both MSE and bias.
7 Empirical Application
//TODO
8 Conclusions
//TODO
Appendix A Functional PIRLS Algorithm details
The maximizer $(\hat{\beta},\hat{f})$ of (4) must satisfy the following system of first-order equations:
$\begin{cases}\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
This system involves derivatives with respect to both parameters $\beta$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}$ is given by:
$\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)x_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (13)
For the second equation of (12), we note that $m(u,\hat{f})$ is the derivative of $m(\hat{f},\hat{f})$. We now compute the terms involving $\mathcal{L}$ only:
$\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right).$
Thus, we obtain:
$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (14)
Define the following quantities:
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
$r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f).$
Thus, (13) can be rewritten as:
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\beta,f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f)x_{jk}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\beta,f)x_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\beta,f)}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\beta,f)x_{ik}l_{i}(\beta,f)$
$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\beta,f)l_{i}(\beta,f))x_{ik},$
where $l_{i}(\beta,f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)}$. Continuing, we have:
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\theta_{i}(\beta,f)x_{ik}$
$=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}.$
Similarly, the Gâteaux derivative of $\mathcal{L}(\beta,f)$ in the direction $u$ is:
$\frac{\partial\mathcal{L}(\beta,u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}.$
Thus, solving (12) is equivalent to finding $(\hat{\beta},\hat{f})$ that satisfies:
$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
Assuming $l_{i}$ and $v_{i}$ are constant, solving (15) is equivalent to finding the minimizer of the functional:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$. Assuming such quantities are constant suggests an iterative computation scheme.
Let $\theta^{(k)}$ be an estimate of $\theta(\beta,f)$ after $k$ iterations. We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\beta^{(k)},f^{(k)})$:
$\theta(\beta,f)\approx\underbrace{\exp(X\beta^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\beta,f)}{\partial\beta}(\beta-\beta^{(k)})+\frac{\partial\theta(\beta,f)}{\partial f}(f-f^{(k)}).$
The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\beta$:
$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\partial\theta_{i}^{(k)}x_{ij},$
which in matrix form is:
$\frac{\partial\theta(\beta,f)}{\partial\beta}=\Theta^{(k)}X,$
where $\Theta^{(k)}=\text{diag}(\theta^{(k)})$.
Similarly, for the derivative with respect to $f$, we have:
$\frac{\partial\theta_{i}(\beta^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\beta^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
Hence, the first-order approximation of $\mathcal{J}_{\mu}(\beta,f)$ in the neighborhood of $(\beta^{(k)},f^{(k)})$ is:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\delta/l-\left(\theta^{(k)}+\Theta^{(k)}X(\beta-\beta^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}-X\beta-f\right)\right\|^{2}+\mu m(f,f).$
Defining $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$, we can rewrite:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X\beta-f)\right\|^{2}+\mu m(f,f).$ (16)
Since ${W}^{(k)}$ is positive definite, $\tilde{\mathcal{J}}_{\lambda}$ is a quadratic form with a unique minimum.

### v2

@@
 Spatial Proportional Hazards Model with Differential Regularization
-Spatial Proportional Hazards Model with Differential Regularization
-Lorenzo Tedesco
+Abstract
+This paper presents a semiparametric proportional hazards model designed to handle spatially varying covariate functions, applicable to both geostatistical and areal data observed on irregular spatial domains.
+The model is estimated through the maximization of a penalized partial likelihood, with a roughness penalty term based on a differential of the spatial field over the target domain.
+The finite element method is employed for efficient estimation, enabling a piecewise polynomial surface representation of the spatial field.
+We apply this method to analyze response time data from the London Fire Brigade.
+Keywords: spatial proportional hazards model, differential regularization, semiparametric model, geostatistical data, areal data, finite element method, spatial field, surivival analysis
 1 Introduction
 Survival analysis has long been an essential tool across various fields, including medicine, engineering, and social sciences, for analyzing time-to-event data.
-In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards model, first introduced by Cox (1972).
-This model assesses the effect of covariates on the hazard, or risk, of an event occurring.
-In recent years, survival analysis has made significant strides in modeling different forms of dependence in survival data.
-This paper explores a scenario where the occurrence of an event is associated with its geographic location, introducing a spatial dependence into the data.
-Specifically, an event’s occurrence may be influenced by spatial factors related to the location where the event happens.
-We refer to the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards model.
-Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions and our method is able to deal with both the framework.
-A comprehensive literature review, including Bayesian approaches, can be found in Hanson and Zhou (2014).
-Our interest lies in a frequentist approach to the spatial Proportional Hazards model, for which various models and methods are now available to facilitate inference.
-Examples include the semiparametric estimating equations method proposed by Li and Lin (2006) and the composite likelihood method Paik and Ying (2012).
-The main issue with these proposals is that the dependence structure does not consider the shape of the domain but only the distance among the locations.
-This can be problematic when the boundary of the domain plays a crucial role in the spatial dependence.
-In this paper, we propose a novel approach based on recent developments in differential regularization, as discussed in Wilhelm and Sangalli (2016).
-One of the main features of this approach is to be able to deal with domains having complex shapes, characterized for instance by strong concavities or holes, and where the shape of the domain influences the behavior of the phenomenon.
-We demonstrate that the methodology introduced in Wilhelm and Sangalli (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data.
+In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards (PH) model, first introduced by Cox, (1972).
+The PH model assesses how covariates influence the hazard, or risk, of an event happening at any given moment in time, relative to a baseline risk.
+This model operates under the assumption that the effect of each covariate is proportional and constant over time, meaning that covariates alter the hazard rate by a multiplicative factor, without assuming a specific underlying time distribution for the event.
+In recent years, survival analysis has advanced significantly in modeling various dependencies within survival data.
+Much research has focused on the relationship between censoring and time variables Deresa and Keilegom, (2024), a critical aspect when handling censored data.
+However, other forms of dependence have received comparatively less attention.
+This paper addresses a scenario where the occurrence of an event is linked to its geographic location, thereby introducing spatial dependence into the analysis.
+Specifically, the probability of an event may be influenced by spatial characteristics associated with the event’s location.
+We call the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards (SPH) model.
+Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions, and our method is able to deal with both the framework.
+A comprehensive literature review, including Bayesian approaches, is available in Hanson and Zhou, (2014).
+In this work, however, we focus on a frequentist approach to the SPH model, for which various inference methods have been developed.
+Examples include the semiparametric estimating equations method by Li and Lin, (2006) and the composite likelihood method by Paik and Ying, (2012).
+A significant limitation of these methods is their reliance on a dependence structure that accounts only for the distance between locations, ignoring the domain’s shape.
+This approach can lead to inaccuracies, especially when spatial dependence is influenced by the domain boundary.
+Additionally, most semiparametric SPH models in the literature use a spline estimator to parameterize the baseline hazard function as infinite-dimensional parameters, rather than leaving this component fully unspecified, as in Cox’s traditional partial likelihood approach (Cox,, 1972).
+This additional parameterization requires estimating more parameters, leading to reduced robustness and accuracy in the estimation process.
+In contrast, our method is grounded in the traditional partial likelihood approach, focusing specifically on the key multiplicative component of the model.
+Unlike other methodologies that introduce spatial dependence through a frailty-based approach—where the spatial domain is clustered and each cluster has a distinct effect on the hazard following a specified distribution—we incorporate a non-parametric spatial field defined over the entire domain to capture spatial dependence in the data.
+For the estimation, we propose a novel method that leverages recent advancements in differential regularization, as discussed by Wilhelm and Sangalli, (2016).
+A key advantage of our approach is its capacity to accommodate domains with complex shapes, such as those characterized by strong concavities or holes, where the domain geometry significantly impacts the behavior of the observed phenomenon.
+We demonstrate that the methodology introduced in Wilhelm and Sangalli, (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data.
 Specifically, we optimize a penalized log-likelihood function that includes a roughness penalty term based on a differential quantity of the spatial field, calculated over the domain of interest.
-This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli, 2016).
-Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field, particularly in domains with complex geometries, where triangulation effectively captures the intricacies of the area.
+This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli,, 2016).
+Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field.
 The paper is organized as follows: In Section 2, we introduce the Spatial Proportional Hazards model.
@@
 We consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with a boundary $\partial\Omega$ that is regular and belongs to $C^{2}(\mathbb{R}^{2})$.
-Let $p_{1},\ldots,p_{n}\in\Omega$ denote $n$ points, where $p_{i}=(p_{1i},p_{2i})$.
-Each point $p_{i}$ is associated with a time-to-event random variable $T_{i}$.
+Let $\bm{p}_{1},\ldots,\bm{p}_{n}\in\Omega$ denote $n$ points, where $\bm{p}_{i}=(p_{1i},p_{2i})$.
+Each point $\bm{p}_{i}$ is associated with a time-to-event random variable $T_{i}$.
 Due to censoring, $T_{i}$ cannot be observed directly; instead, we obtain a realization of $(Y_{i},\Delta_{i})$, denoted by $(y_{i},\delta_{i})$, where $Y_{i}=\min(T_{i},C_{i})$, $\Delta_{i}=I(T_{i}\leq C_{i})$, $C_{i}$ is a censoring random variable, and $I(\cdot)$ is the indicator function.
-Additionally, we observe a covariate vector $x_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$.
-We assume that $T_{i}$ is independent of $C_{i}$ conditional on $x_{i}$.
+Additionally, we observe a covariate vector $\bm{x}_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$.
+We assume that $T_{i}$ is independent of $C_{i}$ conditional on $\bm{x}_{i}$.
 We use a proportional hazard model for the random variables $\{T_{i}\}_{i=1}^{n}$.
-Specifically, let $\lambda_{i}(t\mid x)$ denote the conditional hazard function of $T_{i}$ given $x_{i}=x$.
+Specifically, let $\lambda_{i}(t\mid\bm{x})$ denote the conditional hazard function of $T_{i}$ given $\bm{x}_{i}=\bm{x}$.
 We assume that
-$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(x^{\top}\beta+f(p_{i})),$ (1)
-where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function over $\Omega$ belonging to a suitable functional space $\mathcal{F}$.
-Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $p^{*}\in\Omega$ it holds $f(p^{*})=0$.
-We propose to estimate the regression coefficients $\beta\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
+$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(\bm{x}^{\top}\bm{\beta}+f(\bm{p}_{i})),$ (1)
+where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function from $\Omega$ to $\mathbb{R}$ belonging to a suitable functional space $\mathcal{F}$.
+Thus, we will interpret the model in terms of hazards: the vector of coefficients $\bm{\beta}$ quantifies the influence of the covariates $\bm{x}_{i}$ on the hazard function, so each component of $\beta$ can be interpreted as the log hazard ratio associated with a one-unit increase in the corresponding covariate, holding other factors constant.
+The function $f(\bm{p}_{i})$, instead, captures the spatial variation in hazards that is not accounted for by the covariates.
+Specifically, $f(\bm{p}_{i})$ allows for the identification of localized risk factors affecting the hazard at different points in the domain, revealing areas with higher or lower risks relative to the baseline hazard $\lambda_{0}(t)$.
+Thus, the interpretation of hazards is twofold: it incorporates covariate effects via $\bm{\beta}$ and captures residual spatial variation through $f(\bm{p})$, providing a comprehensive view of risk that addresses both individual and spatial heterogeneity.
+Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $\bm{p}^{*}\in\Omega$ it holds $f(\bm{p}^{*})=0$.
+Note that alternative identifiability conditions can be considered, such as $\int_{\Omega}f=1$.
+We propose to estimate the regression coefficients $\bm{\beta}\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
 […]
-where $\mathcal{L}(\beta,f)$ is the standard partial likelihood defined as
+where $\mathcal{L}(\bm{\beta},f)$ is the standard log-partial-likelihood for the PH model, defined as
 $\mathcal{L}(\beta,f)=\sum_{i:\delta_{i}=1}\left(x_{i}^{\top}\beta+f(p_{i})-\log\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\beta+f(p_{j})\right)\right).$ (3)
@@
 A higher $\mu$ controls the wiggliness of $f$, while a smaller $\mu$ allows for more flexibility in $f$.
-We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al.
-(2013).
-The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\beta}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$.
-As in Wilhelm and Sangalli (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al.
-(1986) for maximizing a penalized log-likelihood functional in generalized additive models.
+We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al., (2013).
+The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\bm{\beta}}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$.
+As in Wilhelm and Sangalli, (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al., (1986) for maximizing a penalized log-likelihood functional in generalized additive models.
 This is because the penalized log-likelihood in (2) involves a function parameter, the spatial field $f$, and the maximization occurs over the space $\mathbb{R}^{q}\times\mathcal{F}$, where $\mathcal{F}$ is infinite-dimensional.
-Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli (2016) to find an approximate solution to the estimation problem.
+Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli, (2016) to find an approximate solution to the estimation problem.
 3 Functional version of the PIRLS algorithm
 We consider the more general penalized log-likelihood function given by:
-$\mathcal{L}_{p}(\beta,f)=\mathcal{L}(\beta,f)-\frac{\mu}{2}m(f,f),$ (4)
+$\mathcal{L}_{p}(\bm{\beta},f)=\mathcal{L}(\bm{\beta},f)-\frac{\mu}{2}m(f,f),$ (4)
 where $\mathcal{L}$ denotes the log-likelihood and $m(\cdot,\cdot):\mathcal{F}\times\mathcal{F}\to\mathbb{R}$ is a bilinear, symmetric, and semi-positive definite form.
@@
 For that, define the following quantities:
-$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
+$\theta_{i}(\bm{\beta},f)$ $=\exp(\bm{x}_{i}^{\top}\bm{\beta}+f(p_{i})),$
 $r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
-$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f),$
-$l_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)},$
-$v_{i}(\beta,f)$ $=\theta_{i}(\beta,f)/l_{i}(\beta,f).$
-First, we show that the problem of maximizing 4 is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
-$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
-where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant.
-As, in fact, $l$ and $V$ also depend on $(\beta,f)$, this suggests an iterative scheme for the solution of the estimation problem.
-Let $(\beta^{(k)},f_{n}^{(k)})$ be the value of $(\beta,f)$ after $k$ iterations of the algorithm, and let $\theta^{(k)}$ the estimate of the vector $\theta$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\theta^{(k)})$.
-Define also $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X^{\top}\beta^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$.
-The first order approximation of $\theta(\beta,f)$ in the neighborhood of the current value of $\theta^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
-$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X^{\top}\beta-f)\right\|^{2}+\mu m(f,f).$ (5)
+$s_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f),$
+$l_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\bm{\beta},f)},$
+$v_{i}(\bm{\beta},f)$ $=\theta_{i}(\bm{\beta},f)/l_{i}(\bm{\beta},f).$
+First, we show that the problem of maximizing (4) is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
+$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\bm{\delta}/\bm{l}-\theta)\|^{2}+\mu m(f,f),$
+where $\bm{\delta}=(\delta_{1},\dots,\delta_{n})^{\top}$, $\bm{l}=(l_{1},\dots,l_{n})^{\top}$, $\bm{\delta}/\bm{l}=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$, and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant.
+As, in fact, $\bm{l}$ and $V$ also depend on $(\bm{\beta},f)$, this suggests an iterative scheme for the solution of the estimation problem.
+Let $(\bm{\beta}^{(k)},f_{n}^{(k)})$ be the value of $(\bm{\beta},f)$ after $k$ iterations of the algorithm, and let $\bm{\theta}^{(k)}$ the estimate of the vector $\bm{\theta}=(\theta_{1},\dots,\theta_{n})^{\top}$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\bm{\theta}^{(k)})$.
+Let $X$ be the $q\times n$ matrix where the column $i$ is equal to $x_{i}$ and define also $\bm{z}^{(k)}=(\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\bm{\theta}^{(k)})+X^{\top}\bm{\beta}^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\bm{\beta}^{(k)},f^{(k)})\theta_{i}^{2}(\bm{\beta}^{(k)},f^{(k)})$.
+The first order approximation of $\theta(\bm{\beta},f)$ in the neighborhood of the current value of $\bm{\theta}^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
+$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\bm{\beta},f)=\left\|({W}^{(k)})^{1/2}(\bm{z}^{(k)}-X^{\top}\bm{\beta}-f)\right\|^{2}+\mu m(f,f).$ (5)
 We can consider the following iterative scheme:
   1. 1.
-    Compute $z^{(k)}$ and $W^{(k)}$ based on $(\beta^{(k)},f_{n}^{(k)})$.
+    Compute $\bm{z}^{(k)}$ and $W^{(k)}$ based on $(\bm{\beta}^{(k)},f_{n}^{(k)})$.
   2. 2.
-    Next, find $\beta$ and $f_{n}$ that minimize (5).
+    Next, find $\bm{\beta}$ and $f_{n}$ that minimize (5).
   3. 3.
-    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\beta}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
+    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\bm{\beta}}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
   4. 4.
-    Define $\beta^{(k+1)}=\hat{\beta}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
+    Define $\bm{\beta}^{(k+1)}=\hat{\bm{\beta}}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
 Here $f_{n}^{(k)}(p^{*})$ is the estimation of $f$ at step $k$ at $p^{*}$.
@@
 […]
-$\tilde{J}_{\lambda}(\beta,f)=\|W^{1/2}(z-X\beta-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
+$\tilde{J}_{\lambda}(\bm{\beta},f)=\|W^{1/2}(z-X^{\top}\bm{\beta}-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
 For simplicity, we omit the dependence on the iteration counter $k$ in this notation and throughout the following sections.
@@
 Typically, constraints can be set on the value of $f$ at the boundary $\partial\Omega$, i.e., $f|_{\partial\Omega}=\gamma_{D}$ (Dirichlet boundary conditions), or on the flux of the function through the boundary, i.e., $\partial_{n}f|_{\partial\Omega}=(\nabla f)^{\top}\mathbf{n}=\gamma_{N}$ (Neumann boundary conditions), where $\mathbf{n}$ is the outward-pointing normal unit vector to the boundary, and $\nabla f=(\partial f/\partial p_{1},\partial f/\partial p_{2})^{\top}$ represents the gradient of $f$.
-If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are termed homogeneous.
+If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are said to be homogeneous.
 It is also possible to apply different boundary conditions to various parts of the boundary, forming a partition of $\partial\Omega$.
@@
 $\mathcal{F}=H^{2}_{n}=\left\{f\in H^{2}\mid(\nabla f)^{\top}\mathbf{n}=0\text{on }\partial\Omega\right\}.$
-For more details on general boundary conditions, see Azzimonti et al. (2014).
+For more details on general boundary conditions, see Azzimonti et al., (2014).
 4.1 Characterization of the Solution to the Penalized Partial Likelihood Problem
@@
 Define $H=X(X^{\top}WX)^{-1}X^{\top}W$ and $Q=I-H$, where $I$ is an identity matrix of the appropriate size.
-For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $u_{n}=(u(p_{1}),\ldots,u(p_{n}))$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{p_{i}\}_{i=1}^{n}$.
-Let $\tilde{\beta}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\beta,f)$ in (4).
-Under these conditions, the following Proposition describes the minimizers $\tilde{\beta}$ and $\tilde{f}$ of the functional (6).
+For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $\bm{u}_{n}=(u(\bm{p}_{1}),\ldots,u(\bm{p}_{n}))^{\top}$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{\bm{p}_{i}\}_{i=1}^{n}$.
+Let $\tilde{\bm{\beta}}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\bm{\beta},f)$ in (4).
+Under these conditions, the following proposition describes the minimizers $\tilde{\bm{\beta}}$ and $\tilde{f}$ of the functional (6).
 Proposition 1.
-There exists a unique pair $(\tilde{\beta},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
-  • $\tilde{\beta}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{f}_{n}=(\tilde{f}(p_{1}),\ldots,\tilde{f}(p_{n}))^{\top}$,
+There exists a unique pair $(\tilde{\bm{\beta}},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
+  • $\tilde{\bm{\beta}}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{\bm{f}}_{n}=(\tilde{f}(\bm{p}_{1}),\ldots,\tilde{f}(\bm{p}_{n}))^{\top}$,
   • $\tilde{f}$ satisfies:
-    $u_{n}^{\top}Q\tilde{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=u_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
+    $\bm{u}_{n}^{\top}Q\bm{\tilde{}}{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=\bm{u}_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
 Using Proposition 1 and the functional version of the PIRLS algorithm from Section 3, we obtain a characterization of the maximum penalized log-likelihood within the functional space $H^{2}_{n}(\Omega)$.
@@
 $H^{1}_{n}=\left\{f\in H^{1}\mid(\nabla f)^{\top}\mathbf{n}=0\text{ on }\partial\Omega\right\}.$
-According to Wilhelm and Sangalli (2016); Sangalli et al. (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
-$\begin{cases}u_{n}^{\top}Q\tilde{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=u_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
+According to Wilhelm and Sangalli, (2016); Sangalli et al., (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
+$\begin{cases}\bm{u}_{n}^{\top}Q\bm{\tilde{}}{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=\bm{u}_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
 for each $(u,v)\in H^{1}_{n}\times H^{1}_{n}(\Omega)$.
@@
 4.3 Finite Elements
-The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni, 2009).
+The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni,, 2009).
 To build a finite element space, we begin by dividing the domain of interest $\Omega$ into smaller subdomains.
@@
 Linear finite elements are created using a basis system where each basis function $\psi_{i}$ is associated with a vertex $\xi_{i}$ (for $i=1,\ldots,K$) of the triangulation $\mathcal{T}$.
-Each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta.
+In particular, each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta.
 Figure 1 provides an example of a linear finite element basis function on a planar mesh, highlighting its localized support.
-Let $\psi=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$.
+Let $\bm{\psi}=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$.
 Any function $h$ in the finite element space $\mathcal{F}_{K}$ can be expressed as a linear combination of the basis functions $\psi_{1},\ldots,\psi_{K}$.
-Let $h=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $h$, such that
+Let $\bm{h}=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $\bm{h}$, such that
 $h(\cdot)=\sum_{j=1}^{K}h_{j}\psi_{j}(\cdot)=\mathbf{h}^{\top}\psi(\cdot).$
@@
 As a result, the problem (8) is approximated by its discrete version: find $(\tilde{f},\tilde{h})\in\mathcal{F}_{K}\times\mathcal{F}_{K}$ such that (8) holds for all $(u,v)\in\mathcal{F}_{K}\times\mathcal{F}_{K}$, where the integrals are now computed over the triangulation $\Omega_{\mathcal{T}}$.
-Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\mathbf{p}_{1},\ldots,\mathbf{p}_{n}$,
-$\Psi=\begin{bmatrix}\psi^{\top}(\mathbf{p}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{p}_{n})\end{bmatrix}$ (9)
+Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\bm{p}_{1},\ldots,\bm{p}_{n}$,
+$\Psi=\begin{bmatrix}\psi^{\top}(\bm{p}_{1})\\ \vdots\\ \psi^{\top}(\bm{p}_{n})\end{bmatrix}$ (9)
 and consider the $K\times K$ matrices
-$\mathbf{R}_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad\mathbf{R}_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
+$R_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad R_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
 Using this notation, for functions $\tilde{f},\tilde{h},u,v\in\mathcal{F}_{K}$, the integrals in (8) can be written as:
-$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}\mathbf{R}_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}\mathbf{R}_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}\mathbf{R}_{0}\mathbf{v},$
+$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}R_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}R_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}R_{0}\mathbf{v},$
 where $\tilde{\mathbf{f}},\tilde{\mathbf{h}},\mathbf{u}$, and $\mathbf{v}$ are the coefficient vectors of the basis expansions of $\tilde{f},\tilde{h},u$, and $v$, respectively.
@@
 The discrete form of (8) is given by the system
-$\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\Psi&\lambda\mathbf{R}_{1}\\ \lambda\mathbf{R}_{1}&\lambda\mathbf{R}_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\mathbf{z}\\ 0\end{bmatrix},$ (10)
+$\begin{bmatrix}-\Psi^{\top}Q\Psi&\lambda R_{1}\\ \lambda R_{1}&\lambda R_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}Q\mathbf{z}\\ \bm{0}\end{bmatrix},$ (10)
 which has a unique solution pair $(\tilde{\mathbf{f}},\tilde{\mathbf{h}})$ corresponding to the coefficients of the basis expansions of $\tilde{f}$ and $\tilde{h}$.
-The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $\mathbf{R}_{0}$ and $(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1})$.
-Let $\mathbf{P}=\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1}$ and $\mathbf{S}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}$.
+Proof.
+The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $R_{0}$ and $(\Psi^{\top}Q\Psi+\lambda R_{1}R_{0}^{-1}R_{1})$. ∎
+Let $P=R_{1}R_{0}^{-1}R_{1}$ and $S=\Psi(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q$.
 Using the functional form of the PIRLS algorithm, and leveraging Propositions 1 and 2, we obtain the following expressions for the maximizers $\hat{\bm{\beta}}$ and $\hat{f}$ of the penalized partial likelihood (2):
-$\hat{\bm{\beta}}=(\mathbf{X}^{\top}\mathbf{W}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{W}(\mathbf{I}-\mathbf{S})\mathbf{z},$
+$\hat{\bm{\beta}}=(X^{\top}WX)^{-1}X^{\top}W(I-S)\bm{z},$
 with $\hat{f}$ identified by the vector
-$\hat{\mathbf{f}}=(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z},$
-and the vector of evaluations of $\hat{f}$ at the $n$ data locations given by
-$\hat{\mathbf{f}}_{n}=\Psi\hat{\mathbf{f}}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z}=\mathbf{S}\mathbf{z},$
-where the vector of pseudo-data $\mathbf{z}$, and the matrices $\mathbf{W},\mathbf{Q}$, and $\mathbf{S}$ are those obtained upon convergence of the PIRLS algorithm.
-The positive definite matrix $\mathbf{P}$ represents the discretization of the penalty term in (2) and (7).
+$\hat{\bm{f}}=(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q\mathbf{z},$
+and the vector of evaluations of $\hat{\bm{f}}$ at the $n$ data locations given by
+$\hat{\bm{f}}_{n}=\Psi\hat{\bm{f}}=\Psi(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q\bm{z}=S\bm{z},$
+where the vector of pseudo-data $\bm{z}$, and the matrices $W,Q$, and $S$ are those obtained upon convergence of the PIRLS algorithm.
+The positive definite matrix $P$ represents the discretization of the penalty term in (2) and (7).
 Due to the variational formulation (8) of the estimation problem, this penalty matrix does not require the computation of second-order derivatives.
-Azzimonti et al.
-(2014) demonstrates that in the finite element space used for discretization, $\mathbf{P}$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
+Azzimonti et al., (2014) demonstrates that in the finite element space used for discretization, $P$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
 4.5 Selecting the Parameter $\mu$ Using Cross-Validation
@@
 However, to avoid subjective decision-making, we propose a more objective, numerical method.
-Following the approach of Cygu et al.
-(2021), we can utilize cross-validation to determine the optimal value of $\mu$.
-Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al., 1996).
+Following the approach of Cygu et al., (2021), we can utilize cross-validation to determine the optimal value of $\mu$.
+Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al.,, 1996).
 To compute the CV-PLD for each $\mu$, we employ $k$-fold cross-validation.
@@
 This procedure is repeated $k$ times, each time with a different fold serving as the validation set.
-We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny, 2019).
+We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny,, 2019).
 The CV-PLD is defined as:
 […]
-where $\mathcal{L}(\hat{\beta}_{-k}(\mu))$ represents the log partial likelihood evaluated at $\hat{\beta}_{-k}$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\beta}_{-k}(\mu))$ is the log partial likelihood evaluated at $\hat{\beta}_{-k}$ on the retained data (all but the excluded fold).
-The $\hat{\beta}_{-k}$ values denote the penalized estimates derived from the retained data.
+where $\mathcal{L}(\hat{\bm{\beta}}_{-k}(\mu),\hat{f}_{-k}(\mu))$ represents the log partial likelihood evaluated at $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\bm{\beta}}_{-k}(\mu),\hat{f}_{-k}(\mu))$ is the log partial likelihood evaluated at $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ on the retained data (all but the excluded fold).
+The $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ values denote the penalized estimates derived from the retained data.
 The optimal value of $\mu$ is the one that minimizes (11).
@@
 This metric, known as the cross-validated C-index, calculates the probability that, for a random pair of individuals, the predicted survival times are correctly ordered according to their actual survival times.
-This method is similar to the implementation available in the survival package (Therneau et al., 2015).
+This method is similar to the implementation available in the survival package (Therneau et al.,, 2015).
 5 Model Version for Areal data
-//TODO
 6 Simulation Results
-To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay (2002); Wood et al.
-(2008).
+To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay, (2002); Wood et al., (2008).
 We use the spatial field $f(x,y)$, depicted in Figure 2(a) and defined as follows:
 $f(x,y)=\begin{cases}-\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y-r)^{2}\leq(r-r_{0})^{2}\text{ and }x\geq 3,\\ -\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,r_{0}\leq y\leq 2r-r_{0},\\ -\frac{1}{8}\left(\arctan\left(-\frac{y}{x}\right)r+\left(\sqrt{x^{2}+y^{2}}-r\right)^{2}+10\right)+c&\text{if }r_{0}^{2}\leq x^{2}+y^{2}\leq(2r-r_{0})^{2}\text{ and }x<0,\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,-(2r-r_{0})\leq y\leq-r_{0},\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y+r)^{2}\leq(r-r_{0})^{2}\text{ and }x>3,\\ 0&\text{otherwise},\end{cases}$
-where $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$, $r=0.5$ and $r_{0}=0.1$.
-We randomly sample $n=100,200,300$ points, denoted by $\{p_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b).
-Figure 2(c) shows the corresponding Delaunay triangulation.
-Each sampled point is associated with an observation $x_{i}$ of a random variable $X_{i}$, which follows a standard normal distribution.
+where $r=0.5$ and $r_{0}=0.1$, and $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$.
+Figure]2(a) shows the values of the spatial field over the horseshoe.
+6.1 Geostatistical Data
+We randomly sample $n=100,200,300$ points, denoted by $\{\bm{p}_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b).
+Figure 2(c) shows the corresponding Delaunay triangulation (Shewchuk,, 2002).
+Each sampled point is associated with an observation $x_{i}$ of a unidimensional random variable $X_{i}$, for which we omit the bold notation, and that follows a standard normal distribution.
 Additionally, we observe $(y_{i},\delta_{i})$, where $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=I(T_{i}\leq C_{i})$.
 The variable $T_{i}$ given $X_{i}=x$ follows a proportional hazards model with baseline hazard $\lambda_{0}(t)=1$ and $\beta_{0}=0.2$, as defined in equation (1), with the form $\lambda(t|x)=\lambda_{0}(t)\exp(x\beta_{0}+f(p_{i}))$.
+This implies that the random variable $T_{i}$ at location ±$\pm p_{i}$ follows an exponential distribution with rate (inverse of the mean) $\exp(x\beta_{0}+f(p_{i}))$.
 To introduce censoring, we set $C_{i}=1$ for each $i$, resulting in approximately 20% censored observations.
-We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ effects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$.
-The finite elements method is applied using a basis of order $1$ constructed on the Toggle the table of contents Delaunay triangulation over the sampled locations.
-(a) Spatial field $f(x,y)$ over the horseshoe domain
-(b) Sampled points over the horseshoe domain
-(c) Mesh based on the sampled points
+We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ affects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$.
+The finite elements method is applied using a basis of order $1$ constructed on the Delaunay triangulation over the sampled locations.
+We conduct $N=250$ simulations for each combination of $n$ and $\mu$.
+An example of the estimated spatial field is given in Figure 2(d).
+(a) Spatial field $f(\bm{p})$ over the horseshoe domain
+(b) Example of sample points over the horseshoe domain for $n=300$.
+(c) Mesh based on the sampled points of Figure 2(b).
+(d) Example of estimation of the spatial field based on the sampled points of Figure 2(b) and setting $\mu=10$.
 Figure 2: Figures illustrating various aspects of the spatial field and sampling.
-We conduct $N=100$ simulations for each combination of $n$ and $\mu$.
-And example of the estimated spatial field in Figure 3.
-This figure compares the true values (Figure 3(a)) with the estimated values (Figure 3(b)), and reports the pointwise error.
-In table
-(a) True Values
-(b) Estimated Values
-(c) Error Values
-Figure 3: Comparison of true values, estimated values, and error values.
-$n$ $\mu$ MSE of $f$ (sd) Bias of $\beta$ (sd)
-100 0.5 0.168 (0.112) -0.008 (0.130)
-1 0.168 (0.118) -0.007 (0.130)
-3 0.177 (0.140) -0.006 (0.130)
-5 0.186 (0.165) -0.006 (0.130)
-10 0.190 (0.188) -0.007 (0.129)
-200 0.5 0.157 (0.119) 0.008 (0.096)
-1 0.160 (0.129) 0.007 (0.096)
-3 0.168 (0.150) 0.007 (0.095)
-5 0.180 (0.169) 0.007 (0.094)
-10 0.185 (0.179) 0.008 (0.093)
-300 0.5 0.141 (0.105) -0.003 (0.077)
-1 0.142 (0.115) -0.002 (0.077)
-3 0.143 (0.129) 0.000 (0.077)
-5 0.144 (0.136) 0.000 (0.077)
-10 0.148 (0.149) 0.000 (0.077)
-Table 1: Simulation results for varying sample sizes ($n$) and smoothing parameters ($\mu$), showing the mean squared error (MSE) of the spatial field estimator $f$ and the bias of the parameter $\beta$.
-Standard deviations (sd) are reported in parentheses for both MSE and bias.
+We report in Table 1 the simulation results that highlights how changes in the smoothing parameter $\mu$, censoring levels, and sample sizes affect key performance metrics.
+the bias values for $\beta$ remain close to zero across most scenarios, indicating unbiased estimates, with slight variations at higher levels of $\mu$.
+The standard deviation (Sd) decreases as the sample size $n$ increases, suggesting that larger samples improve estimate stability.
+Additionally, the 95% confidence interval coverage probability (CP95) stays close to the nominal level, particularly in larger samples, supporting the reliability of interval estimates.
+For the spatial field $f$, we reported the Mean Squared Error (MSE)defined as $\text{MSE}=n^{-1}\sum_{i=1}^{n}(\hat{f}(\bm{p}_{i})-f(\bm{p}_{i}))^{2}$, as a performance metric.
+we observe that the MSE values generally decrease as $n$ grows, suggesting improved precision with more data.
+Higher censoring percentages (e.g., 40%) introduce a slight increase in MSE and its standard deviation (Sd MSE), indicating reduced accuracy under increased censoring.
+Regression coefficient $\beta$ Spatial field $f$
+$\mu$ Cens. % $n$ bias Sd CP95 MSE Sd MSE
+0.5 20% 100 -0.001 0.121 0.944 0.177 0.125
+200 -0.010 0.078 0.964 0.155 0.111
+40% 100 -0.002 0.145 0.944 0.227 0.222
+200 -0.013 0.091 0.960 0.150 0.096
+1 20% 100 -0.001 0.122 0.944 0.172 0.126
+200 -0.010 0.077 0.968 0.157 0.121
+40% 100 -0.003 0.145 0.932 0.206 0.171
+200 -0.012 0.091 0.960 0.153 0.110
+3 20% 100 -0.001 0.121 0.948 0.177 0.146
+200 -0.009 0.077 0.972 0.158 0.137
+40% 100 -0.003 0.145 0.936 0.206 0.178
+200 -0.012 0.091 0.964 0.155 0.128
+5 20% 100 -0.001 0.121 0.948 0.185 0.164
+200 -0.009 0.077 0.972 0.161 0.146
+40% 100 -0.003 0.145 0.940 0.216 0.196
+200 -0.011 0.090 0.964 0.159 0.141
+10 20% 100 -0.002 0.121 0.948 0.196 0.177
+200 -0.008 0.077 0.972 0.166 0.163
+40% 100 -0.002 0.145 0.940 0.229 0.217
+200 -0.010 0.090 0.964 0.163 0.157
+Table 1: Simulation results for estimating the regression coefficient $\beta$ and spatial field $f$ across different values of the smoothing parameter $\mu$, varying censoring percentages (Cens.
+%) and sample sizes $n$.
+The performance metrics include bias, standard deviation (Sd), and 95% confidence interval coverage probability (CP95) for $\beta$, as well as mean squared error (MSE) and standard deviation of MSE (Sd MSE) for $f$.
+The results demonstrate the impact of $\mu$, censoring level, and sample size on the precision and accuracy of the estimates.
+6.2 Areal Data
+//TODO
 7 Empirical Application
-//TODO
+In this section, we apply the proposed method to analyze the London Fire Brigade (LFB) response time data.
+The dataset is available in the R package spatsurv (Taylor and Rowlingson,, 2017) and is further explored in Taylor, (2017).
+The original data can also be accessed via the London Datastore Fire and Authority, (2015).
+Our analysis focuses on data from 2009, specifically on fires occurring in residential dwellings, as these are associated with the highest number of fatalities.
+The response variable of interest is the time taken for the first fire engine to arrive at the scene of a dwelling fire.
+We excluded observations where the response time was missing, and the resulting dataset contains no censored data.
+In Figure 3, we display the observed locations of these incidents, highlighting the spatial distribution of fires across London.
+The blue dots represent the locations of the observed fires, while the grey boundary outlines the geographical area under study.
+As can be seen from the image, the area excludes the River Thames, which corresponds to a natural boundary that could potentially influence the arrival time of the first fire engine.
+Given the irregular path of the river, it was necessary to smooth the boundary of the domain.
+After smoothing the domain and excluding points that fall into the river or outside the region of interest, the dataset results in 6657 observed locations, though less than 0.01% of the points were omitted.
+Some observations had identical locations, and while different strategies could be used, such as jittering the points, we chose to randomly omit the duplicated locations.
+This led to a dataset of 5949 unique locations, resulting in the removal of approximately 10% of the data points, which still leaves a sufficiently large dataset for analysis.
+Figure 3: Observed fire locations in London.
+The blue points represent the geographical locations of fire incidents in residential dwellings across the city.
+The red border outlines the area under study, excluding the River Thames as a natural boundary.
+Figure 4: Distribution of observed response times of the first fire engine to reach fire incidents. The color intensity represents the density of the response times across different locations.
+Figure 5: Fire station intensity map.
+The blue intensity highlights the estimated proximity of fire stations to the observed fire locations, with red contours indicating areas of higher station density.
+It is reasonable to expect that the response time to a fire service call is influenced by the proximity of the nearest fire station.
+However, in practice, the closest station may already be responding to another incident.
+As explored in Taylor and Rowlingson, (2017), we include a scaled fire station intensity measure as a covariate in the model, which accounts for the proximity of fire stations to fire locations.
+A mesh was constructed based on the observed locations using first-order basis functions.
+To determine the smooth parameter $\mu\in\{1,4,7,\dots,100\}$, we applied the CV-PLD method discussed in Section 4.5, utilizing a 10-fold cross-validation approach.
+The optimal $\mu$ was found to be 46.
+The estimation algorithm terminates if the Euclidean norm difference between consecutive estimates of the spatial field is less than 0.01, with convergence typically occurring within 5 iterations.
+The estimated regression coefficient for the fire station density covariate is $\beta=0.0667$, with a standard deviation of $1.41\times 10^{-5}$, indicating a statistically significant effect at the 1% level.
+In Figure 6, we present the estimated spatial field.
+The figure highlights how the river acts as a natural boundary, constraining the fire engines’ ability to respond promptly to incidents on the opposite side of its banks.
+This observation suggests that geographical features significantly impact response efficiency and should be carefully considered when interpreting spatial variations in the data.
+In Figure 7, we provide a cross-section of the spatial field, highlighted as a vertical line in Figure 6.
+This cross-section shows the estimated value of the spatial field at longitude 535000 as a function of latitude.
+We observe a distinct change point around latitude 180000, corresponding to the river’s location, where the hazard level immediately drops by approximately 80%.
+It is arguable that the river effect might be artificially induced by omitting the river space from the domain and that the presence of bridges along the boundary effectively mitigated issues related to water.
+While the choice of domain should ideally be guided by a priori knowledge about the specific problem, it would nonetheless be valuable to conduct a sensitivity analysis regarding domain selection.
+Although this analysis is beyond the scope of this paper, we plan to address it in future studies.
+Figure 6: Estimated spatial field of fire engine response.
+The red contour outlines areas where geographical boundaries, particularly the River Thames, influence the response efficiency of fire engines.
+The vertical dotted line corresponds to the section considered in Figure 7
+Figure 7: Estimated spatial field at Latitude 535000, where it is possible to observe a change point at Latitude 180000 corresponding to the presence of the river.
 8 Conclusions
-//TODO
+This paper introduces a novel SPH model that integrates differential regularization and the finite element method to effectively address spatial dependencies in survival data observed over irregular domains.
+Our methodology expands on the traditional semiparametric proportional hazards model by incorporating a non-parametric spatial field component to account for spatial dependence, providing a flexible approach to capture unobserved geographical heterogeneity in hazard rates across a spatial domain.
+The SPH model framework offers several advantages over existing methods.
+First, our approach employs differential regularization, which controls the smoothness of the estimated spatial field, allowing for adaptation to complex spatial boundaries.
+Secondly, by applying the finite element method on triangulated spatial meshes, we achieve computationally efficient and accurate estimates, particularly in irregularly shaped regions where spatial dependencies do not follow straightforward geometric patterns.
+The proposed model demonstrated promising results in both simulation and empirical studies.
+Simulations confirmed that the model yields accurate estimates of both the covariate effects and the spatial component across various sample sizes and levels of censoring.
+In future work, we aim to enhance the model’s adaptability by investigating alternative penalty structures, including anisotropic differential penalties, to more accurately capture directional dependencies in specific spatial datasets.
+Additionally, further research will focus on the asymptotic properties of the estimators to strengthen the model’s theoretical foundations.
+Sensitivity analysis concerning the choice of the boundary domain will also be pursued.
+Finally, we plan to extend the model to incorporate time-varying covariates within the spatial proportional hazards framework, enabling application to more dynamic spatial survival contexts, such as real-time risk prediction and event-driven spatial analysis.
 Appendix A Functional PIRLS Algorithm details
-The maximizer $(\hat{\beta},\hat{f})$ of (4) must satisfy the following system of first-order equations:
-$\begin{cases}\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
-This system involves derivatives with respect to both parameters $\beta$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
-For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}$ is given by:
-$\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)x_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (13)
+The maximizer $(\hat{\bm{\beta}},\hat{f})$ of (4) must satisfy the following system of first-order equations:
+$\begin{cases}\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
+This system involves derivatives with respect to both parameters $\bm{\beta}$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
+For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}$ is given by:
+$\frac{\partial\mathcal{L}(\hat{\bm{\beta}},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)\bm{x}_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right).$ (13)
 For the second equation of (12), we note that $m(u,\hat{f})$ is the derivative of $m(\hat{f},\hat{f})$. We now compute the terms involving $\mathcal{L}$ only:
-$\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
-$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
-$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
-$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right).$
+$\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right)$
+$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right)$
+$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
+$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}+o(t)\right).$
 Thus, we obtain:
-$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (14)
+$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\bm{\beta}},\hat{f}+\tau u)-\mathcal{L}(\hat{\bm{\beta}},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(\bm{x}_{j}^{\top}\hat{\bm{\beta}}+\hat{f}(p_{j})\right)}\right).$ (14)
 Define the following quantities:
-$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
+$\theta_{i}(\bm{\beta},f)$ $=\exp(\bm{x}_{i}^{\top}\bm{\beta}+f(p_{i})),$
 $r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
-$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f).$
+$s_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f).$
 Thus, (13) can be rewritten as:
-$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\beta,f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f)x_{jk}$
-$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\beta,f)x_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\beta,f)}$
-$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\beta,f)x_{ik}l_{i}(\beta,f)$
-$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\beta,f)l_{i}(\beta,f))x_{ik},$
-where $l_{i}(\beta,f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)}$. Continuing, we have:
-$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\theta_{i}(\beta,f)x_{ik}$
-$=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}.$
+$\frac{\partial\mathcal{L}(\bm{\beta},f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\bm{\beta},f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f)\bm{x}_{jk}$
+$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\bm{\beta},f)\bm{x}_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\bm{\beta},f)}$
+$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\bm{\beta},f)x_{ik}l_{i}(\bm{\beta},f)$
+$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\bm{\beta},f)l_{i}(\bm{\beta},f))\bm{x}_{ik},$
+where $l_{i}(\bm{\beta},f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\bm{\beta},f)}$. Continuing, we have:
+$\frac{\partial\mathcal{L}(\bm{\beta},f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\bm{\beta},f)^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{\theta_{i}(\bm{\beta},f)l_{i}(\beta,f)^{-1}}\theta_{i}(\bm{\beta},f)x_{ik}$
+$=\sum_{i=1}^{n}\frac{l_{i}(\bm{\beta},f)^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{\theta_{i}(\bm{\beta},f)l_{i}(\bm{\beta},f)^{-1}}\frac{\partial\theta_{i}(\bm{\beta},f)}{\partial\beta_{k}}.$
 Similarly, the Gâteaux derivative of $\mathcal{L}(\beta,f)$ in the direction $u$ is:
-$\frac{\partial\mathcal{L}(\beta,u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}.$
-Thus, solving (12) is equivalent to finding $(\hat{\beta},\hat{f})$ that satisfies:
-$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
+$\frac{\partial\mathcal{L}(\bm{\beta},u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},u)}{\partial f}.$
+Thus, solving (12) is equivalent to finding $(\hat{\bm{\beta}},\hat{f})$ that satisfies:
+$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\bm{\beta},f)}{v_{i}}\frac{\partial\theta_{i}(\bm{\beta},u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
 Assuming $l_{i}$ and $v_{i}$ are constant, solving (15) is equivalent to finding the minimizer of the functional:
-$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
-where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$. Assuming such quantities are constant suggests an iterative computation scheme.
-Let $\theta^{(k)}$ be an estimate of $\theta(\beta,f)$ after $k$ iterations. We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\beta^{(k)},f^{(k)})$:
-$\theta(\beta,f)\approx\underbrace{\exp(X\beta^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\beta,f)}{\partial\beta}(\beta-\beta^{(k)})+\frac{\partial\theta(\beta,f)}{\partial f}(f-f^{(k)}).$
-The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\beta$:
-$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\partial\theta_{i}^{(k)}x_{ij},$
+$\mathcal{J}_{\mu}(\bm{\beta},f)=\|V^{-1/2}(\bm{\delta}/\bm{l}-\theta)\|^{2}+\mu m(f,f),$
+where $\bm{\delta}/\bm{l}=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$, $\bm{v}=(v_{1},\dots,v_{n})^{\top}$, $V=\text{diag}(\bm{v})$ and $diag$ denotes a diagonal matrix with the $\bm{v}$ elements along its main diagonal, and zeros elsewhere.
+Assuming such quantities are constant suggests an iterative computation scheme.
+Let $\bm{\theta}^{(k)}$ be an estimate of $\theta(\bm{\beta},f)$ after $k$ iterations.
+We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\bm{\beta}^{(k)},f^{(k)})$:
+$\theta(\bm{\beta},f)\approx\underbrace{\exp(X\bm{\beta}^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\bm{\beta},f)}{\partial\beta}(\bm{\beta}-\bm{\beta}^{(k)})+\frac{\partial\theta(\bm{\beta},f)}{\partial f}(f-f^{(k)}).$
+The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\bm{\beta}$:
+$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\theta_{i}^{(k)}x_{ij},$
 which in matrix form is:
-$\frac{\partial\theta(\beta,f)}{\partial\beta}=\Theta^{(k)}X,$
-where $\Theta^{(k)}=\text{diag}(\theta^{(k)})$.
+$\frac{\partial\bm{\theta}(\bm{\beta},f)}{\partial\bm{\beta}}=\Theta^{(k)}X,$
+where $\Theta^{(k)}=\text{diag}(\bm{\theta}^{(k)})$.
 Similarly, for the derivative with respect to $f$, we have:
-$\frac{\partial\theta_{i}(\beta^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\beta^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
-Hence, the first-order approximation of $\mathcal{J}_{\mu}(\beta,f)$ in the neighborhood of $(\beta^{(k)},f^{(k)})$ is:
-$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\delta/l-\left(\theta^{(k)}+\Theta^{(k)}X(\beta-\beta^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
-$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}-X\beta-f\right)\right\|^{2}+\mu m(f,f).$
-Defining $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$, we can rewrite:
-$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X\beta-f)\right\|^{2}+\mu m(f,f).$ (16)
+$\frac{\partial\theta_{i}(\bm{\beta}^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\bm{\beta}^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
+Hence, the first-order approximation of $\mathcal{J}_{\mu}(\bm{\beta},f)$ in the neighborhood of $(\bm{\beta}^{(k)},f^{(k)})$ is:
+$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\bm{\delta}/\bm{l}-\left(\bm{\theta}^{(k)}+\Theta^{(k)}X(\bm{\beta}-\bm{\beta}^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
+$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\bm{\theta}^{(k)})+X\bm{\beta}^{(k)}+f^{(k)}-X\bm{\beta}-f\right)\right\|^{2}+\mu m(f,f).$
+Defining $\bm{z}^{(k)}=(\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\theta^{(k)})+X\bm{\beta}^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\bm{\beta}^{(k)},f^{(k)})\theta_{i}^{2}(\bm{\beta}^{(k)},f^{(k)})$, we can rewrite:
+$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\bm{\beta},f)=\left\|({W}^{(k)})^{1/2}(\bm{z}^{(k)}-X\bm{\beta}-f)\right\|^{2}+\mu m(f,f).$ (16)
 Since ${W}^{(k)}$ is positive definite, $\tilde{\mathcal{J}}_{\lambda}$ is a quadratic form with a unique minimum.
+Appendix B Proof of Proposition 1
+The proof of Proposition 1 follows the approach outlined in Appendix C of Sangalli et al., (2013). We begin by recalling the Lax-Milgram theorem (Quarteroni and Quarteroni,, 2009).
+Theorem 1 (Lax-Milgram).
+Let $\mathcal{F}$ be a Hilbert space, $G(\cdot,\cdot):\mathcal{F}\times\mathcal{F}\to\mathbb{R}$ a continuous and coercive bilinear form, and $F:\mathcal{F}\to\mathbb{R}$ a linear and continuous functional.
+Then, there exists a unique solution to the following problem:
+$\text{find }u\in\mathcal{F}\text{ such that }G(u,v)=F(v),\quad\forall v\in\mathcal{F}.$
+Moreover, if $G(\cdot,\cdot)$ is symmetric, then $u\in\mathcal{F}$ is the unique minimizer in $\mathcal{F}$ of the functional $B:\mathcal{F}\to\mathbb{R}$, defined as
+$B(u)=G(u,u)-2F(u).$
+We also require the following result.
+Lemma 1.
+The bilinear and symmetric form $G:H^{2}_{n0}(\Omega)\times H^{2}_{n0}(\Omega)\to\mathbb{R}$, defined as
+$G(v,u)=\bm{u}_{n}^{\top}Q\bm{v}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta v),$
+is continuous and coercive.
+Proof.
+We recall the definitions of the norm $\|\cdot\|_{H^{2}}$ and the semi-norm $|\cdot|_{H^{2}}$:
+$\|u\|_{H^{2}}=\sum_{|\alpha|\leq 2}\|\partial^{\alpha}u\|_{L^{2}},\quad|u|_{H^{2}}=\|\Delta u\|_{L^{2}},\quad\forall u\in H^{2}.$
+First, note that the semi-norm $|\cdot|_{H^{2}}$ and the norm $\|\cdot\|_{H^{2}}$ are equivalent in $H^{2}_{n0}(\Omega)$ (Quarteroni and Quarteroni,, 2009), i.e., there exists $C_{0}>0$ such that $|u|_{H^{2}(\Omega)}\geq C_{0}\|u\|_{H^{2}(\Omega)},\,\forall u\in H^{2}_{n0}(\Omega)$.
+Then, we have:
+$G(u,u)=\bm{u}_{n}^{\top}Q\bm{u}_{n}+\lambda\int_{\Omega}(\Delta u)^{2}\geq\lambda\int_{\Omega}(\Delta u)^{2}=\lambda|u|_{H^{2}(\Omega)}^{2}\geq\lambda C_{0}\|u\|_{H^{2}(\Omega)}^{2};$
+thus, $G(\cdot,\cdot)$ is coercive.
+We now show the continuity of $G(\cdot,\cdot)$.
+Since $H^{2}(\Omega)\subset C^{0}(\Omega)$ and since the norms $\|\cdot\|_{\infty}$ and $\|\cdot\|_{2}$ are equivalent on $\mathbb{R}^{n}$, there exists a constant $C_{1}$ such that $\|v_{n}\|_{\infty}\leq C_{1}\|v\|_{H^{2}(\Omega)},\,\forall v\in H^{2}(\Omega)$.
+Since $Q$ is symmetric, its largest eigenvalue $\rho$ is non-negative.
+Then we have:
+$G(u,v)=\bm{u}_{n}^{\top}Q\bm{v}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta v)\leq\rho\|\bm{u}_{n}\|_{\infty}\|\bm{v}_{n}\|_{\infty}+\lambda|u|_{H^{2}(\Omega)}|v|_{H^{2}(\Omega)}\leq\max\{\rho C_{1}^{2},\lambda C_{0}^{2}\}\|u\|_{H^{2}(\Omega)}\|v\|_{H^{2}(\Omega)}.$
+Hence, the bilinear form $G(\cdot,\cdot)$ is also continuous. ∎
+We can now provide the proof of Proposition 1.
+Proof.
+First, given $f\in H^{2}_{n0}$, the unique minimizer of the functional $\tilde{J}_{\lambda}(\beta,f)$ is given by:
+$\tilde{\bm{\beta}}(f)=(X^{\top}WX)^{-1}X^{\top}W(\bm{z}-\bm{f}_{n}).$
+To show this, we take the derivative of $\tilde{J}_{\lambda}(\bm{\beta},f)$ with respect to $\bm{\beta}$:
+$\frac{\partial\tilde{J}_{\lambda}(\bm{\beta},f)}{\partial\bm{\beta}}=-2X^{\top}W(\bm{z}-\bm{f}_{n})+(X^{\top}WX)\bm{\beta}.$
+Since $X$ is a full-rank matrix and $W$ is invertible (each diagonal entry of $W$ is strictly positive by construction), $X^{\top}WX$ is also invertible.
+Finally, the necessary condition $\frac{\partial\tilde{J}_{\lambda}(\tilde{\bm{\beta}},f)}{\partial\bm{\beta}}=0$ is satisfied if and only if $\tilde{\bm{\beta}}$ is given by the above expression.
+Since for fixed $f$, $\tilde{J}_{\lambda}(\bm{\beta},f)$ is clearly convex, $\tilde{\bm{\beta}}$ is indeed a minimum.
+Substituting $\tilde{\bm{\beta}}$ into the objective function, we obtain the following form of the functional:
+$\tilde{J}_{\lambda}(f)=\bm{z}^{\top}Q\bm{z}-2\bm{f}_{n}^{\top}Q\bm{z}+\bm{f}_{n}^{\top}Q\bm{f}_{n}+\lambda\int_{\Omega}(\Delta f)^{2}.$
+To optimize this functional with respect to $f$ alone, the problem reduces to finding $\tilde{f}\in H^{2}_{n0}$ that minimizes:
+$J_{\lambda}^{*}(f)=\bm{f}_{n}^{\top}Q\bm{f}_{n}+\lambda\int_{\Omega}(\Delta f)^{2}-2\bm{f}_{n}^{\top}Q\bm{z}.$
+We can express $J_{\lambda}^{*}(f)=G(f,f)-2F(f)$, where
+$G(u,v)=\bm{u}_{n}^{\top}Q\bm{v}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta v)\quad\text{and}\quad F(v)=v_{n}^{\top}Q\bm{z}.$
+Lemma 1 ensures that $G(\cdot,\cdot)$ is a coercive and continuous bilinear form on $H^{2}_{n0}\times H^{2}_{n0}$; moreover, $F$ is trivially a linear and continuous functional on $H^{2}_{n0}$.
+By applying the Lax-Milgram Theorem and leveraging the symmetry of $G(\cdot,\cdot)$, the minimizer of the functional $J_{\lambda}^{*}$ is the function $\tilde{f}\in H^{2}_{n0}$ such that
+$\bm{u}_{n}^{\top}Q\tilde{\bm{f}}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=u_{n}^{\top}Q\bm{z},\quad\forall u\in H^{2}_{n0}.$
+Thus, we conclude that $\tilde{f}$ exists and is unique, and consequently, $\tilde{\bm{\beta}}$ also exists and is unique. ∎